cathrinerose23:

What quantity of pepper would you need to make a triple recipe? Greek salad INGREDIENTS: 7 1 2 cups lettuce 1 cup red onions 1 cup olives 1 1 4 cups cucumber 1 cup tomatoes 1 1 2 cups feta cheese 1 3 8 teaspoons salt 3 7 8 teaspoons pepper Write your answer as a fraction or as a whole or mixed number.

cathrinerose23:

oka so im guesing that i gotta add the numbers right?

Hero:

Do you know how to screenshot the question?

cathrinerose23:

nah not on this computer

Hero:

Having some trouble interpreting what you posted above.

umm:

I can try and see if I can put it together if you'd like, Hero.

Hero:

Good luck.

cathrinerose23:

go head and try

Hero:

You're able to copy and paste but not screenshot? Seems strange. What kind of computer is it? What operating system?

umm:

What quantity of pepper would you need to make a triple recipe? Greek salad INGREDIENTS: 7 1/2 cups lettuce 1 cup red onions 1 cup olives 1 1/4 cups cucumber 1 cup tomatoes 11/2 cups feta cheese 13/8 teaspoons salt 3 7/8 teaspoons pepper Write your answer as a fraction or as a whole or mixed number. I believe is given?

cathrinerose23:

yes thx u

cathrinerose23:

i couldnt get the whole thing on there

Hero:

Okay so what is your question regarding this problem @cathrinerose23

cathrinerose23:

@Hero i gotta find what quantity of pepper would you need to make a triple recipe? and Write your answer as a fraction or as a whole or mixed number.

Hero:

Is not the quantity already listed in the recipe @cathrinerose23 ?

cathrinerose23:

its gotta be enough for a triple recipe

cathrinerose23:

idk these questions on this website are dumb as ever

Hero:

Okay, so take whatever the quantity is and multiply it times 3

cathrinerose23:

oka

Hero:

In other words, the quantity is $$3 + 7/8$$ If you multiply that by three it is $$3(3 + 7/8)$$

cathrinerose23:

oka

Hero:

So you have to apply the distributive property here.

cathrinerose23:

oka so 3 plus 7/8 is 3.875 aint it

Hero:

It is, however, it is best to apply the distributive property first: \(a(b + c) = ab + ac\)

Hero:

Try to keep the whole numbers separate from the fractions for now.

cathrinerose23:

and then 3(3+7/8) is 11.625

cathrinerose23:

oka

Hero:

It's best to learn how to do it without a calculator. Otherwise you might run into a problem if you ever have to take a test that forbids calculators

cathrinerose23:

oh yea

cathrinerose23:

so now what do i do?

Hero:

The key here is that you have to write your answer as a whole number and a fraction. They don't want the answer in decimals.

Hero:

So in other words, you have to do it by hand.

cathrinerose23:

oka

Hero:

Apply the distributive property: \(3(3 + 7/8) = 3*3 + 3*(7/8)\)

Hero:

So @cathrinerose23 what is 3*3?

cathrinerose23:

9

Hero:

Very good and what is 3*(7/8)? Hint multiply the 3 and the 7 and leave the denominator the same.

cathrinerose23:

21/8

Hero:

Very good. so now we have \(3 + 21/8\)

Hero:

Now we need to convert the improper fraction to a proper fraction

cathrinerose23:

ok

Hero:

The easiest way to do that is to express the 21 as the HIGHEST multiple of 8 plus the remainder.

Hero:

Any ideas?

cathrinerose23:

well 3 = 21/8 is 5.625 and that converted into a fraction is 45/8

cathrinerose23:

from what i know of

Hero:

Well no, what we want to do is express the answer as a whole number + proper fraction.

Hero:

45/8 is still an improper fraction

cathrinerose23:

oka

Hero:

So, what are some multiples of 8. Can you name one?

cathrinerose23:

ummm... u got 8 duh, 16, 24, 32 u no what im gettin at right

Hero:

Right we're only looking for the highest multiple of 8 that doesn't exceed 21

Hero:

Do you know which number that is?

cathrinerose23:

8*3

cathrinerose23:

i mean 8*2 which is 16

Hero:

Exactly so we can express the result as 3 + (16 + 5)/8

cathrinerose23:

yea

Hero:

But we're still not done just yet

Hero:

In proper form this looks like \(3 + \dfrac{16 + 5}{8}\)

Hero:

Next we split the fraction to two fractions. Do you know how to do that?

cathrinerose23:

ok

cathrinerose23:

now what

cathrinerose23:

what we probably have to do is solve and convert and i got 23/8 and that converted is 2 7/8

Hero:

No, that's not it. Follow the steps I was showing you.

Hero:

What did you get after splitting the fraction to two fractions?

cathrinerose23:

splitting the two fractions?

cathrinerose23:

the one that had the 16+5?

Hero:

yes that one.

Hero:

How do you split \(\dfrac{16 + 5}{8}\) to two fractions?

Hero:

By the way the rule for splitting a fraction to two fractions is \(\dfrac{a + b}{c} = \dfrac{a}{c} + \dfrac{b}{c}\)

cathrinerose23:

wouldnt it be 16/8 and 5/8

Hero:

Yes correct so altogether what we have is \(3 + \dfrac{16}{8} + \dfrac{5}{8}\). From here, do you see anything that can be simplified?

cathrinerose23:

🤷‍♂️🤷‍♂️🤷‍♂️

Hero:

What is 16 divided by 8?

cathrinerose23:

2

Hero:

I'm hoping you did not use a calculator for that.

cathrinerose23:

no

cathrinerose23:

thats easy math

Hero:

Okay, so now we have \(3 + 2 + \dfrac{5}{8}\). Do you see a way to simplify that?

cathrinerose23:

i may be slow at math but im not dumb lol🤣🤣😁😁

cathrinerose23:

yes

Hero:

So what is the final result?

cathrinerose23:

whole number right

Hero:

A whole number PLUS a fraction.

cathrinerose23:

3+2 is 5

cathrinerose23:

would i add that to 5/8?

Hero:

Correct.

cathrinerose23:

oka but wouldnt that be a decimal?

Hero:

No, you don't write it as a decimal. You simply express it as a whole number plus a fraction

cathrinerose23:

so 6/25?

Hero:

Absolutely not.

cathrinerose23:

hmmm im dumb as ever

Hero:

You express it as \(a + \dfrac{b}{c}\).

cathrinerose23:

😒😒😒😢😢😫😫

cathrinerose23:

6 2/5

Hero:

\(a\) represents the whole number and \(\dfrac{b}{c}\) represents the fraction.

cathrinerose23:

or not those #s

Hero:

No, not that either go back to the numbers we were dealing with before you started talking about \(6\) and \(25\).

Hero:

There were two things we were adding together. What were they?

cathrinerose23:

5 and 5/8

cathrinerose23:

??

Hero:

Bingo

cathrinerose23:

ok

cathrinerose23:

thats what the answer was this whole freaking time??!!!

cathrinerose23:

OMFGGGGGGGG

Hero:

But I avoid doing it that way because you can get it confused with multiplication.

cathrinerose23:

im so damn slow at this subject im not passing the sats

Hero:

You're in HS?

Hero:

\(5 + \dfrac{5}{8}\) sorry is the answer

cathrinerose23:

oh so i was close right

Hero:

You had it the whole time. Just takes some getting used to. That's all it is. Mathematics is just something you already know. You just need to familiarize yourself with it more.

cathrinerose23:

ohhhh

cathrinerose23:

why cant i realize that??

Hero:

You're realizing it now

cathrinerose23:

i am